OpenStudy (deannamarx01):

Celia is starring at the clock waiting for school to end so that she can go to track practice. She notices the 4-inch-long minute hand is rotating around the clock and marking off time like degrees on a unit circle. Part 1: How many radians does the minute hand move from 1:25 to 1:50? (hint find the number of degrees per minute first) Part 2: How far does the tip of the minute hand travel during that time? Show all work please. :( @bmk614

OpenStudy (anonymous):

do you're own work

OpenStudy (ask_if_im_a_ghost):

Part 1: To find out how many degrees per minute the minute hand turns, you divide the number of degrees the minute hand turns by however many minutes it took. So to give you a hint; in an hour, how many degrees does a minute hand turn?

OpenStudy (bmk614):

2pi is the total radians for a circle. 60 minutes per hour. 2pi/60=pi/30 50-25=25 minutes 25*pi/30=5pi/6

OpenStudy (anonymous):

part 2??? @bmk614

OpenStudy (ask_if_im_a_ghost):

Part 2: The distance traveled is the radius times how many radians it has covered. So multiply the length of the minute hand by how many radians it traveled.

OpenStudy (bmk614):

5pi/6*4 10pi/3

OpenStudy (ask_if_im_a_ghost):

Correct, the minute hand travels 10*Pi/3 inches